how many liters of water must be added to 10 liters of a 95% acid solution to make an 80% acid solution?
Answer by ptaylor(2198)

Let x=amount of water that must be added
Now we know that the amount of pure acid before the water is added (0.95*10) has to equal the amount of pure acid after the water is added(0.80(10+x)). Sooooo
0.80(10+x)=0.95*10
8+0.80x=9.5
0.80x=1.5
x=1.875 liters=1 7/8 liter---amount of water that needs to be added
Check
0.95*10=0.80*11.875
9.5=9.5
Hope this helps---ptaylor